OpenStudy (anonymous):

The volume of a rectangular prism is 105 cubic centimeters. The length is 7 centimeters and the height is 5 centimeters. What is the width?

Vocaloid (vocaloid):

volume = length*width*height 105 = 7*width*5 solve for width

OpenStudy (anonymous):

so I just multiply 105*7*8???????????

Vocaloid (vocaloid):

no... you need to divide both sides of the equation by 7 and 5

OpenStudy (anonymous):

V=l*w*h so 105=7*w*8 solve for'' w''

Vocaloid (vocaloid):

where is 8 coming from? it's nowhere in the problem

OpenStudy (anonymous):

woah my bad haha 5 not 8